I have tried this problem four times---My book says the answer should be -13/2
I keep getting different answers----What is the step by step procedure?
PROBLEM: 2/5x-5=12/5x+8

Answer by suresh(20)

2/5x - 5 = 12/5x + 8
Subtract 12/5x on both sides,
2/5x - 12/5x - 5 = 8
-10/5x - 5 = 8
Add 5 on both sides,
-10/5x = 8+5 = 13
Taking reciprocal,
-5x/10 = 1/13
x = - 10 / 5(13)=-2 / 13.
So your answer is x=-2/13.